OpenStudy (anonymous):

square root of 2x+6 minus square root of x+10 equals 1 x=15 is this right?

OpenStudy (anonymous):

2(15)+6= 30+6 36

OpenStudy (anonymous):

are you saying 15 is the answer that you got

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\sqrt{2x+6} - \sqrt{x+10}=1\]

OpenStudy (anonymous):

is this your equation

OpenStudy (anonymous):

yes

OpenStudy (anonymous):